OpenStudy (anonymous):

The length of a rectangular frame is 3 cm more than the width. The area inside the frame is 40 square cm. Find the width of the frame.

OpenStudy (anonymous):

l = 3+w a = lw a = w(3+w) = 40 40 = w(3+w) w=5

OpenStudy (anonymous):

Area is LxW You can make the width equal to w and the length will be w+3 And the rest seemed to have appeared above :-)